OpenStudy (anonymous):

Estimate the area under the curve f(x) = 16 – x^2 from x = 0 to x = 3 by using three inscribed (under the curve) rectangles. Answer to the nearest integer.

OpenStudy (anonymous):

Would I take the integral of the function from 0 to 3?

OpenStudy (dinamix):

yup this it @Ephemera use integral u know how use it right ?

OpenStudy (anonymous):

\[\int\limits_{0}^{3}16-x ^{2} \] That would be it correct?

OpenStudy (dinamix):

yup

OpenStudy (dinamix):

\[\left[ 16x-\frac{ x^3 }{ 3 } \right]\]

OpenStudy (dinamix):

(16(3)- 3^3/3) - 0 = 48-9 = 39 this the area @Ephemera

OpenStudy (anonymous):

@dinamix just solved it myself and got 39 as well. Appreciated.

OpenStudy (anonymous):

@dinamix I got this one wrong too....could it be a faulty system?

OpenStudy (dinamix):

so what fault u sopke @Ephemera

OpenStudy (anonymous):

What?